OpenStudy (anonymous):

The circumference of a circle is 50.24 centimeters. Find the radius of the circle. (Use 3.14 for π.)

OpenStudy (anonymous):

circumference of circle \[= \Pi r ^{2}\] you are given the circumference , pi is given so only one unknown parameter you can calculate \[r = \sqrt{\frac{ c }{ \Pi }} \] where c is the circumference I took only the +ve vale for the square root as it is radius "length" so it can't be a negative value

OpenStudy (anonymous):

its one of these ansers

OpenStudy (anonymous):

8 cm 16 cm 4 cm 2 cm

OpenStudy (anonymous):

it will be \[\sqrt{\frac{ 50.24 }{ 3 }} = 4 cm\]

OpenStudy (anonymous):

sorry 3.14

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

urw :)